# The Energy Loss Phenomenon in RLHF: A New Perspective on Mitigating Reward Hacking

### v2

@@ Appendix I Human Evaluation in Reward Hacking Identification
In this section, we present some qualitative examples comparison in Figures 19, 20, and 21. We observe that due to constraints directly applied to the LLM’s output space, such as KL divergence and response length, PPO w/ KL and PPO w/ LP suffer from limited policy exploration space, resulting in restricted performance gains in RLHF. Moreover, as PPO w/ LP addresses only length bias, it remains susceptible to reward hacking (e.g., overly cautious responses) to some extent. In contrast, by solely constraining the energy loss in the LLM’s final layer, our EPPO effectively mitigates reward hacking while enabling a broader policy optimization space, leading to superior RLHF performance.

### v3

@@ Appendix I Human Evaluation in Reward Hacking Identification
In this section, we present some qualitative examples comparison in Figures 20, 21, and 22. We observe that due to constraints directly applied to the LLM’s output space, such as KL divergence and response length, PPO w/ KL and PPO w/ LP suffer from limited policy exploration space, resulting in restricted performance gains in RLHF. Moreover, as PPO w/ LP addresses only length bias, it remains susceptible to reward hacking (e.g., overly cautious responses) to some extent. In contrast, by solely constraining the energy loss in the LLM’s final layer, our EPPO effectively mitigates reward hacking while enabling a broader policy optimization space, leading to superior RLHF performance.